# Equivariant Symmetry-Aware Head Pose Estimation for Fetal MRI

### v4

@@ 5.4 Results
We first evaluate performance on in-domain research data, and then assess generalization to target clinical domains. Table 1 reports the performance statistics on the test volumes from each dataset, and Fig. 3 shows example alignments. Since FireANTs and Fetal-Align do not involve network training beyond our segmentation network, which is only trained on Research-Fetal, we exclude these baseline methods from our training experiments on dHCP.
@@
E(3)-Pose accurately predicts pose and significantly outperforms other direct pose regression methods, showcasing the advantage of modeling rotational equivariance. The relative rankings of 3DPose-Net, 6DRep, and RbR highlight the stability gains of estimating continuous and higher-dimensional rotation parametrizations [zhou2019]. E(3)-Pose takes advantage of this concept by regressing a continuous, overcomplete output. On dHCP test data, E(3)-Pose is outperformed by template-based methods. This strategy is able to effectively exploit the information advantage of a template volume coming from the same subject when the intensity distributions of both volumes are highly detailed and similar, eliminating alignment ambiguities. Furthermore, the 1.8s inference time of FireANTs (cf. 0.3s for E(3)-Pose) is prohibitive for real-time deployment. Finally, Fetal-Align also yields accurate pose estimates, demonstrating the reliability of landmark detection methods in volumes where landmarks are clearly visible.
@@
E(3)-Pose yields accurate pose estimates in clinical acquisitions of younger ages, outperforming all baseline methods. This difference is more pronounced when training on dHCP, which consists of higher-SNR volumes and presents an even larger distribution shift from clinical data compared to Research-Fetal. E(3)-Pose is the only method that maintains its performance in both training settings, highlighting the robust generalization advantages of our method. For further insight, we study the performance as a function of GA. Fig. 4 reveals that template-based methods (FireANTs and EquiTrack) break down on younger fetuses, where underdeveloped, spherical brain shapes and the lack of structure differentiation within the brain [zhan2013] lead to alignment ambiguities (Fig. 3). 6DRep and RbR remain relatively stable across GAs when trained on Research-Fetal: by removing dependence on specific anatomical markers, direct pose regression is less sensitive to the lack of fine-grained details. However, pose regression with standard CNNs struggles to generalize to younger GAs when trained on dHCP, suggesting that these methods are prone to overfitting to the training distribution. In contrast, E(3)-Pose provides accurate results across GAs and training datasets.
@@
E(3)-Pose achieves accurate pose estimation in navigator volumes and significantly outperforms all baselines (Table 1). While all baseline methods struggle with the challenging task of generalizing from dHCP to Navigators, E(3)-Pose remains robust to the substantial domain gap. Importantly, E(3)-Pose continues to maintain accuracy in the most challenging cases, which occur when the effects of low resolution, low SNR, and artifact obstruction combine to eliminate most asymmetric cues (Fig. 3, columns 4-6). Nevertheless, our method successfully recovers head pose, even in volumes with high uncertainty. In contrast, template registration methods are prone to errors when spin history artifacts disrupt the similarity between input and template intensity distributions, and Fetal-Align remains highly sensitive to eye visibility, which is often limited by artifact obstruction or blurring due to voxel size. Moreover, E(3)-Pose generally yields more stable estimates compared to CNN-based pose regression, highlighting that enforcing pose equivariance and anatomical symmetry on the network parameters is an effective regularization technique that enables generalization to out-of-distribution clinical volumes and stability under pose ambiguity.
Table 2: Ablation study. Mean $\pm$ standard deviation statistics for rotation error () and average absolute error (AAD, $m$). Best score is in bold. * indicates statistical significance compared to E(3)-Pose (p<0.05p\!<\!0.05, pairwise Wilcoxon test).
@@ 5.5 Ablation study
To study performance on challenging out-of-distribution cases, we evaluate ablations on clinical datasets (Table 2). Appendix I provides results on research datasets.
@@
We ablate the modeling of rotational pose equivariance by replacing all E(3)-equivariant convolutions with standard convolutions equivariant to only translations. The results show that rotation equivariance is critical for generalization to clinical data, especially when the domain gap between training and testing data is significant. This outcome is in line with prior work [lawrence2022, sannai2019, bulusu2022] and our earlier comparison of E(3)-Pose to non-equivariant RbR, 6DRep, and 3DPose-Net.
@@
We now examine the effect of two alternative loss functions that are not invariant under left-right reflections. We first train our network by replacing the pseudovector loss term in Eq. (9) with that of the standard vectors. Next, we replace our objective with the widely used geodesic loss [salehi2019]. Table 2 shows that both loss functions significantly reduce performance in Navigators, emphasizing the importance of combining both symmetry-invariant architectures and loss functions for stable performance under high ambiguity. Lastly, ablating the simulated spin history artifact simulations during training shows that this augmentation critically enables generalization to Navigators, where volumes contain real spin history artifacts.
@@ 6 Discussion and Conclusion
E(3)-Pose does not explicitly model motion history or identify high-uncertainty cases, both of which could help to mitigate errors in noisy clinical volumes. Thus, future work will investigate equivariant approaches for estimating probability distributions over SO(3), conditioned on motion trajectories [klee2023, rangaprasad2016]. Beyond fetal imaging, it is interesting to investigate applications of our method when approximate reflectional and/or rigid symmetries [wang2022] are present, such as cardiac, lung, and adult brain imaging. Lastly, we expect that our method could be adapted to objects with more complex (e.g., $n$-fold rotations) symmetries (Appendix K).
@@
Supplementary Material
Appendix A Automated slice prescription in fetal brain MRI
[…]
In this section, we describe our clinical application in further detail.
@@ Appendix A Automated slice prescription in fetal brain MRI
Radiological assessment of abnormal fetal brain development relies on volumetric fetal brain MRI, which provides greater contrast and resolution than standard ultrasound examinations. Due to sensitivity to fetal motion, current clinical practice does not involve direct acquisition of 3D MRI scans [glenn2009]. Instead, stacks of 2D MRI slices are acquired in the three target anatomical orientations (i.e., sagittal, coronal, and axial). Slices that closely align with the target anatomical orientation are critical for abnormality detection. For instance, non-oblique midline sagittal slices are necessary for detecting agenesis of the corpus callosum [gholipour2014]. Furthermore, inter-slice fetal motion is often correlated with poor spatial coverage. Even small coverage gaps can substantially reduce diagnostic potential. For example, the proportion of the total fetal brain volume constituted by the cerebellum ranges from only 0.03-0.06 in second- and third-trimester fetuses [cai2020]. Adaptive slice prescription based on the fetal head motion thus holds potential to improve the quality of prenatal care.
Slices are T2-weighted and acquired with the half-Fourier acquisition single-shot turbo spin-echo (HASTE) MRI sequence ($1.25\!\times\!1.25$$m$ pixels, 3mm slice thickness, $320\!\times\!320$mm FOV, TR=2.5s, TE=106ms, GRAPPA R=2, Partial Fourier=5/8, $90^{\circ}$ flip angle). Fig. 6 shows examples of diagnostic 2D MRI slices in the target anatomical orientations, as well as oblique slices caused by fetal motion. To automatically prescribe the acquisition plane of each 2D slice according to the current fetal head pose, we rapidly acquire a 3D navigator volume before each slice. Navigator volumes are acquired with the echo-planar imaging (EPI) sequence (4-6mm isotropic voxels, $324\!\times\!324\!\times\!120$mm FOV, TR=29ms, TE=14ms, 5 flip angle). Low-resolution, low-energy navigator volumes are not of diagnostic quality but suffice for head pose estimation.
@@
We aim to rapidly estimate the 6-DoF head pose in each navigator volume and use it to adjust the imaging plane of the next slice (Fig. 7). Formally, at the start of the sequence, we define the target imaging planes $P_{1},...,P_{K}\!\in\!\text{SE(3)}$ of all $K$ slices in the stack, relative to the canonical fetal head frame. The rotational and translational components of $P_{k}$ indicate the target anatomical orientation and the relative position of the slice within the stack, respectively. Before the $k$th slice in the stack, we acquire a navigator volume, from which we rapidly estimate the head pose $T_{k}=t_{k}\circ R_{k}$ and prescribe the next slice imaging plane to be $\tilde{P}_{k}=T_{k}P_{k}$. We also translate the field-of-view (FOV) of the next navigator volume around the fetal head based on $t_{k}$. We include this step to ensure full head capture for accurate pose estimation. Because the time interval between every navigator volume and the next slice is 1s, we require the runtime of pose estimation to be less than 1s.
Figure 8: Real-time feedback loop system. Our implementation runs on a 3T Siemens scanner, connected to a server hosted on a GPU-enabled laptop. On the server (right), E(3)-Pose separately estimates the rotation $R_{k}$ and translation $t_{k}$ of the fetal head pose $T_{k}$, based on the navigator volume $f_{k}$ sent by the scanner (left). We apply translation $t_{k}$ to center the next navigator volume FOV around the fetal head. Future work will apply the full rigid transform $T_{k}$ to prescribe the next slice to account for the current head pose. Finally, we send the prescription parameters to the scanner for the next acquisition.
In-utero implementation
We developed a real-time feedback loop system using existing software [wighton2024] that runs this sequence on a 3T Siemens scanner (Fig. 8). Specifically, we have implemented the translation estimation step of E(3)-Pose and used it to translate the FOV of every navigator volume on pregnant patients in real time. The videos in the linked project page demonstrate how the navigator volumes are automatically shifted to follow the translational movements of the fetal head. Future work will implement real-time prescription of slices, using the full pose $T_{k}$ returned by E(3)-Pose in every navigator volume.
@@ Appendix B E(3)-CNN implementation details
E(3)-CNNs necessitate special poolings and non-linearities, since standard versions only satisfy Eq. (2) for zero-order fields. Specifically, higher order (l>0l\!>\!0) fields require (i) norm-based pooling, and (ii) special non-linearities learned by the network as multiplicative scalar fields: $f(x)\sigma(\kappa_{l,p,0,0}\!\ast\!f(x))$, where $\sigma$ is the sigmoid function, and $\ast$ is the convolution operator [weiler2018].
@@
We crop input volumes to the E(3)-CNN around the predicted brain mask to $64^{3}$ voxels, scaled such that the brain occupies $60\%$ of the cropped volume size. Our E(3)-CNN architecture has 4 levels, each consisting of 2 E(3)-equivariant convolutions [geiger2022, diaz2024] with kernel size $5\!\times\!5\!\times\!5$, followed by instance normalization and equivariant max-pooling. Following [diaz2024], the first level has 8, 4, and 2 even- and odd-parity output features of orders $l=0,1,2$, respectively, and the feature count is doubled in each level. We use ReLU and tanh activations for even and odd scalar features respectively, and equivariant sigmoid activations for higher-order features. For the radial basis functions $\varphi_{m}$ in the $k^{\text{th}}$ level, we use $\varphi_{m}(r)\!=\!8.433573\text{sus}(x+m-1)\text{sus}(1-m-x)$, where sus is the soft unit step function implemented in [diaz2024] and parametrized by $m\in\{0,r/4,r/2,3r/4,r\}$, where $r=2^{k-2}$.
@@ Appendix C Segmentation network architecture
Following [silva2024], we implement the brain segmentation function $\psi$ with a standard 3D U-Net. Our architecture has 4 levels with 16, 32, 64, and 128 output channels, respectively, each consisting of 2 convolutional layers with kernel size $3\!\times\!3\!\times\!3$, followed by batch normalization and ELU activation [clevert2015]. Inputs are padded to $128^{3}$ voxels. While we only use the predicted brain mask for inference, we train the U-Net to additionally segment the eyes, with the aim of enhancing the overall robustness via multi-task learning [gao2020].
Appendix D Training details
@@
We train the segmentation U-Net for 1000 epochs using Adam optimization [kingma2015] with batch size 4 and learning rate $10^{-4}$. We use a weighted sum of the cross-entropy and Dice losses [milletari2016], with weights of 1 and 0.5, respectively. We re-weight both loss terms by 8, 2, and 1 for the brain, eyes, and background classes, respectively. During training and inference, we resample all input volumes to 3mm isotropic voxels before padding to $128^{3}$ voxels.
We spatially augment training volumes with random rotations, translations, and scaling uniformly sampled from SO(3), $30,30]^{3}$mm, and $[0.5,1.3]$, respectively. We also simulate low resolution in 75% of training volumes, with isotropic voxel size sampled uniformly from [3,8][3,8]$m$. We add Gaussian noise with $\sigma$ sampled uniformly from $[0,0.03]$, random gamma correction with $\log{\gamma}$ sampled uniformly from [0.8,0][-0.8,0], and bias field artifacts. For the bias field simulation, we use existing software [perezgarcia2021] to model the artifact as a linear combination of polynomial basis functions of order 3, with coefficients sampled uniformly from $[0,0.5]$. For spin history artifact simulation, we sample $n_{\text{slice}}$ uniformly from the unit sphere and $\sigma$ uniformly from [1.5,2.3][1.5,2.3]$m$. We sample $c_{\text{slice}}$ such that the likelihood of a point is inversely proportional to its distance from the brain boundary. In navigator volumes, when the artifact intersects regions near the boundary of the brain, there is greater uncertainty in disambiguating the brain from external uterine tissue. Thus, we provide the network with more examples of these challenging cases during training.

### v5

@@ 5.4 Results
We first evaluate performance on in-domain research data, and then assess generalization to target clinical domains. Table 1 reports the performance statistics on the test volumes from each dataset, and Fig. 3 shows example alignments.
@@
E(3)-Pose accurately predicts pose and significantly outperforms other direct pose regression methods, showcasing the advantage of modeling rotational equivariance. E(3)-Pose is outperformed by template-based methods on dHCP test data, where the intensity distribution of the input volume is similar to the template at higher resolution and SNR, eliminating alignment ambiguities. Here, template registration is able to effectively exploit the information advantage of a subject-specific template aligned to the GT pose. The accuracy of Fetal-Align indicates that precise landmark detection supports reliable computation of the head pose, consistent with our landmark-based GT pose definition (Appendix E).
@@
E(3)-Pose yields accurate pose estimates in clinical acquisitions of younger ages, outperforming all baseline methods. This difference is more pronounced when training on dHCP, which consists of higher-SNR volumes and presents an even larger distribution shift from clinical data compared to Research-Fetal. E(3)-Pose is the only method that maintains its performance in both training settings, highlighting the robust generalization advantages of our method. For further insight, we study the performance as a function of GA. Fig. 4 reveals that template-based methods (FireANTs and EquiTrack) break down on younger fetuses, where underdeveloped, spherical brain shapes and the lack of structure differentiation within the brain [79] lead to alignment ambiguities (Fig. 3). 6DRep and RbR remain relatively stable across GAs when trained on Research-Fetal: by removing dependence on specific anatomical markers, direct pose regression is less sensitive to the lack of fine-grained details. However, pose regression with standard CNNs struggles to generalize to younger GAs when trained on dHCP, suggesting that these methods are prone to overfitting to the training distribution. In contrast, E(3)-Pose provides accurate results across GAs and training datasets.
@@
E(3)-Pose achieves accurate pose estimation in navigator volumes and significantly outperforms all baselines (Table 1). While all baseline methods struggle with the challenging task of generalizing from dHCP to Navigators, E(3)-Pose remains robust to the substantial domain gap. Importantly, E(3)-Pose successfully recovers head pose even in volumes with high uncertainty, which occur when the effects of low resolution, low SNR, and artifact obstruction combine to eliminate most asymmetric cues (Fig. 3, columns 4-6). In contrast, template registration methods are prone to errors when spin history artifacts disrupt the similarity between input and template intensity distributions, and Fetal-Align remains highly sensitive to eye visibility, which is often limited by artifact obstruction or blurring due to voxel size. Moreover, E(3)-Pose generally yields more stable estimates compared to CNN-based pose regression, highlighting that enforcing pose equivariance and anatomical symmetry on the network parameters is an effective regularization technique that enables generalization to out-of-distribution clinical volumes and stability under pose ambiguity.
Table 2: Ablation study. Mean $\pm$ standard deviation statistics for rotation error () and average absolute error (AAD, $m$). Best score is in bold. * indicates statistical significance compared to E(3)-Pose at p<0.05p\!<\!0.05 compared to E(3)-Pose (hierarchical permutation test for Navigators, pairwise Wilcoxon for all other datasets). See Tables 7-8 for subject-level statistics in Navigators.
@@ 5.5 Ablation study
To study performance on challenging out-of-distribution cases, we evaluate ablations on clinical datasets (Table 2). Appendix K provides results on research datasets, and Fig. 13 provides example alignments.
@@
We ablate the modeling of rotational pose equivariance by replacing all E(3)-equivariant convolutions with standard convolutions equivariant to only translations. The results show that rotation equivariance is critical for generalization to clinical data, especially when the domain gap between training and testing data is significant. This outcome is in line with prior work [41, 61, 6] and our comparison to non-equivariant RbR, 6DRep, and 3DPose-Net.
@@
We now examine the effect of two alternative loss functions that are not invariant under left-right reflections (Appendix K). We first train our network by replacing the pseudovector loss term in Eq. (7) with that of the standard vectors. Next, we replace our objective with the widely used geodesic loss [60]. Table 2 shows that both loss functions significantly reduce performance in Navigators, emphasizing the importance of combining both symmetry-invariant architectures and loss functions for stable performance under high ambiguity. Lastly, ablating the simulated spin history artifact during training shows that this augmentation critically enables generalization to Navigators, where volumes contain real spin history artifacts.
@@ 6 Discussion and Conclusion
Our symmetry assumption, which is only an approximation based on the low resolution of navigator volumes, may break under large lateralized pathology. In these cases, the three-vector variant of E(3)-Pose, which does not assume left-right symmetry, may be more robust. In addition, E(3)-Pose does not explicitly model motion history or identify high-uncertainty cases, both of which could help to mitigate errors in noisy clinical volumes. Future work will investigate equivariant approaches for estimating probability distributions over SO(3), conditioned on motion trajectories [39, 57]. See Appendix M for further discussion of limitations.
@@
Supplementary Material
Appendix A Automated slice prescription
@@
In this section, we describe our clinical application in detail.
@@
Radiological assessment of abnormal fetal brain development relies on volumetric fetal brain MRI, which provides greater contrast and resolution than standard ultrasound examinations. Due to the high sensitivity of 3D MRI acquisitons to fetal motion, current clinical practice acquires stacks of diagnostic-quality 2D MRI slices in the three target anatomical directions (i.e., sagittal, coronal, and axial) [24]. Slices that closely align to the target anatomical orientation are critical for abnormality detection. For instance, non-oblique midline sagittal slices are necessary for detecting agenesis of the corpus callosum [23]. Furthermore, inter-slice fetal motion is often correlated with poor spatial coverage. Even small coverage gaps can substantially reduce diagnostic potential: for example, in second- and third-trimester fetuses, the proportion of the total brain volume constituted by the cerebellum ranges from only 0.03 to 0.06 [7]. Adaptive slice prescription based on the fetal head motion thus holds potential to improve the quality of prenatal care.
Slices are T2-weighted and acquired with the half-Fourier acquisition single-shot turbo spin-echo (HASTE) MRI sequence ($1.25\!\times\!1.25$$m$ pixels, 3mm slice thickness, $320\!\times\!320$mm FOV, TR=2.5s, TE=106ms, GRAPPA R=2, Partial Fourier=5/8, $90^{\circ}$ flip angle). Fig. 6 shows examples of slices in the target anatomical orientations, as well as oblique slices caused by fetal motion. To automatically prescribe the imaging plane of each slice according to the current fetal head pose, we rapidly acquire a 3D navigator volume before each slice [18]. Navigator volumes are acquired with the echo-planar imaging (EPI) sequence (4-6mm isotropic voxels, $324\!\times\!324\!\times\!120$mm FOV, TR=29ms, TE=14ms, 5 flip angle). Low-resolution, low-energy navigator volumes are not of diagnostic quality but suffice for head pose estimation.
@@
We aim to rapidly estimate the 6-DoF head pose in each navigator volume and use it to adjust the imaging plane of the next slice (Fig. 7). Formally, at the start of the sequence, we define the target imaging planes $P_{1},...,P_{K}\!\in\!\text{SE(3)}$ of all $K$ slices in the stack, relative to the canonical fetal head frame. The rotational and translational components of $P_{k}$ indicate the target anatomical orientation and the relative position of the slice within the stack, respectively. Before the $k$th slice in the stack, we acquire a navigator volume, from which we rapidly estimate the head pose $T_{k}=t_{k}\circ R_{k}$ and prescribe the imaging plane of the next slice to be $\tilde{P}_{k}=T_{k}P_{k}$. We also translate the field-of-view (FOV) of the next navigator volume around the fetal head based on $t_{k}$. We include this step to ensure full head capture for accurate pose estimation. Because the time interval between every navigator and the next slice is 1s, we require the runtime of pose estimation to be less than 1s.
Figure 8: Feedback loop system. Our implementation runs on a 3T Siemens scanner, connected to a server hosted on a GPU-enabled laptop. On the server (right), E(3)-Pose separately estimates the rotation $R_{k}$ and translation $t_{k}$ of the fetal head pose $T_{k}$, based on the navigator volume $f_{k}$ sent by the scanner (left). We apply translation $t_{k}$ to center the next navigator volume FOV around the fetal head. Finally, we send the prescription parameters to the scanner for the next acquisition.
In-utero implementation [54]
We implemented our full automation framework on a GPU-enabled laptop server connected to a 3T Siemens scanner using existing software [73] (Fig. 8). We have performed preliminary evaluations of real-time navigator FOV translation on pregnant volunteers; the videos available at ramyamut.github.io/e3-pose demonstrate how the navigators are automatically shifted to follow the translational movements of the fetal head. Ongoing work is evaluating real-time slice prescription, using the full pose $T_{k}$ returned by E(3)-Pose in every navigator volume.
@@ Appendix B E(3)-CNN implementation details
E(3)-CNNs necessitate special poolings and non-linearities, since standard versions only satisfy Eq. (1) for zero-order fields. Specifically, higher order (l>0l\!>\!0) fields require (i) norm-based pooling, and (ii) special non-linearities learned by the network as multiplicative scalar fields: $f(x)\sigma(\kappa_{l,p,0,0}\!\ast\!f(x))$, where $\sigma$ is the sigmoid function, and $\ast$ is the convolution operator [71].
@@
We crop input volumes to the E(3)-CNN around the predicted brain mask to $64^{3}$ voxels, scaled such that the brain occupies $60\%$ of the cropped volume size. Our E(3)-CNN architecture has 4 levels, each consisting of 2 E(3)-equivariant convolutions [21, 14] with kernel size $5\!\times\!5\!\times\!5$, followed by instance normalization and equivariant max-pooling. Following [14], the first level has 8, 4, and 2 even- and odd-parity output features of orders $l=0,1,2$, respectively, and the feature count is doubled in each level. We use ReLU and tanh activations for even and odd scalar features respectively, and equivariant sigmoid activations for higher-order features. For the radial basis functions $\varphi_{m}$ in the $k^{\text{th}}$ level, we use $\varphi_{m}(r)\!=\!8.433573\text{sus}(x+m-1)\text{sus}(1-m-x)$, where sus is the soft unit step function implemented in [14] and parametrized by $m\in\{0,r/4,r/2,3r/4,r\}$, where $r=2^{k-2}$.
@@ Appendix C Segmentation network architecture
Following [64], we implement the brain segmentation function $\psi$ with a standard 3D U-Net. Our architecture has 4 levels with 16, 32, 64, and 128 output channels, respectively, each consisting of 2 convolutional layers with kernel size $3\!\times\!3\!\times\!3$, followed by batch normalization and ELU activation [10]. Inputs are padded to $128^{3}$ voxels. While we only use the predicted brain mask for inference, we train the U-Net to additionally segment the eyes, with the aim of enhancing the overall robustness via multi-task learning [19].
Appendix D E(3)-CNN loss function
Below we show that in our training objective (Eq. 7), the pseudovector term $|\sin{(\theta_{x}})|$ and the vector terms $|\sin{(\theta_{y}}/2)|$ and $|\sin{(\theta_{z}}/2)|$ are monotonically decreasing functions of the vector dot product $\hat{e}_{k}\cdot e_{k}$ and its square $(\hat{e}_{k}\cdot e_{k})^{2}$, respectively.
$|\sin{(\theta_{x}})|$ $=\sqrt{1-\cos^{2}(\theta_{x})}$ (8)
$=\sqrt{1-(\hat{e}_{k}\cdot e_{k})^{2}}$
$|\sin{(\theta_{\{y,z\}}/2})|$ $=\sqrt{\frac{1-\cos{(\theta_{\{y,z\}})}}{2}}$
$=\sqrt{\frac{1-(\hat{e}_{\{y,z\}}\cdot e_{\{y,z\}})}{2}}$
We note that this objective is invariant under left-right reflections, i.e., $|\sin{(\theta_{x}})|$ remains unchanged under the transformation $e_{x}\!\rightarrow-e_{x}$. We observed experimentally that our choice of loss function produced more robust results than dot product, geodesic loss, and other commonly used variants (Table 2).
Appendix E Pose annotations
Manual
[…]
Semi-automated
To feasibly annotate poses in Navigators, a much larger dataset (1210 total volumes), we utilized algorithmically generated poses to assist with annotation. Specifically, we first algorithmically obtained $P_{k}$ values for all slices, using optimization-based slice-to-volume registration [76] (recall from Appendix A that slices are prescribed according to $\tilde{P}_{k}=T_{k}P_{k}$). Since all $\tilde{P}_{k}$ values were available to us as part of the acquisition parameters, we then computed $T_{k}=\tilde{P}_{k}P_{k}^{-1}$ as the GT pose for every navigator volume preceding the $k$th slice in the stack. Lastly, we manually corrected the poses in each navigator volume.
Appendix F Training details
@@
We train the segmentation U-Net for 1000 epochs using Adam optimization [38] with batch size 4 and learning rate $10^{-4}$. We use a weighted sum of the cross-entropy and Dice losses [51], with weights of 1 and 0.5, respectively. We re-weight both loss terms by 8, 2, and 1 for the brain, eyes, and background classes, respectively. During training and inference, we resample all input volumes to 3mm isotropic voxels before padding to $128^{3}$ voxels.
We spatially augment training volumes with random rotations, translations, and scaling uniformly sampled from SO(3), $30,30]^{3}$mm, and $[0.5,1.3]$, respectively. We also simulate low resolution in 75% of training volumes, with isotropic voxel size sampled uniformly from [3,8][3,8]$m$. We add Gaussian noise with $\sigma$ sampled uniformly from $[0,0.03]$, random gamma correction with $\log{\gamma}$ sampled uniformly from [0.8,0][-0.8,0], and bias field artifacts. For the bias field simulation, we use existing software [55] to model the artifact as a linear combination of polynomial basis functions of order 3, with coefficients sampled uniformly from $[0,0.5]$. For spin history artifact simulation, we sample $n_{\text{slice}}$ uniformly from the unit sphere and $\sigma$ uniformly from [1.5,2.3][1.5,2.3]$m$. We sample $c_{\text{slice}}$ such that the likelihood of a point is inversely proportional to its distance from the brain boundary. In navigator volumes, when the artifact intersects regions near the boundary of the brain, there is greater uncertainty in disambiguating the brain from external uterine tissue. Thus, we provide the network with more examples of these challenging cases during training.
Table 3: Performance statistics for translation estimation. Mean $\pm$ standard deviation for translation error ($m$) are reported. Best score is in bold. * indicates statistical significance at p<0.05p\!<\!0.05 compared to E(3)-Pose (hierarchical permutation test for Navigators, pairwise Wilcoxon for all other datasets). We compare E(3)-Pose only to methods that estimate translation differently from E(3)-Pose. We evaluate the instance of E(3)-Pose trained on Research-Fetal since translation estimation relies only on brain segmentation. FireANTs runtime is 1.8s, prohibiting real-time deployment. See Table 6 for subject-level statistics in Navigators.
Research-Fetal test dHCP test Clinical-Young Navigators
FireANTs [36]   2.0 $\pm$ 2.3* 0.5 $\mathbf{\pm}$ 1.6* 0.6 $\mathbf{\pm}$ 0.5* 4.2 $\pm$ 3.1*
EquiTrack [3]   1.3 $\pm$ 0.9* 1.1 $\pm$ 0.7* 1.6 $\pm$ 1.6* 4.4 $\pm$ 3.2*